How to solve an equation by first clearing the fraction.
5/6x=-1/3

Answer by RagingGod(3)

5/6x =-1/3
Multiply the left hand side and the right hand side by 6.
5/6x X 6 = -1/3 X6
5x = -2
x = -2/5 = -0.4